# Fault-Tolerant Control for System Availability and Continuous Operation in Heavy-Duty Wheeled Mobile Robots

### v1

@@ Abstract
Index Terms:
Robust control, fault-tolerant control, wheeled mobile robot, hydraulic system.
@@ I Introduction
Both light-duty and heavy-duty wheeled mobile robots (HD-WMRs) can perform tasks that are inefficient, unsafe, or unfeasible for humans. However, while light-duty WMRs are typically used in structured environments, like warehouses and laboratories, HD-WMRs excel in challenging applications in rough terrains and under heavy-load conditions [1]. Many HD-WMRs are increasingly equipped with in-wheel hydraulic drive systems, each actuated by its motor and paired with front and rear steering mechanisms to improve maneuverability under heavy loads, enhance responsiveness to off-road conditions, and enable independent power delivery [2].
@@
To address the mentioned challenges, this paper introduces a novel model-free hierarchical control with fault accommodation (MFHCA) framework for highly complex and multi-stage HD-WMRs. It aims to accommodate both sensor and actuator faults passively in hydraulically powered HD-WMRs with independently controlled wheels, while ensuring continuous operation. Contributions of this work are summarized, as follows: 1) a new mathematical representation of the motion dynamics of HD-WMRs, incorporating various sensor and actuator modes, such as healthy, stuck failure, no signal, inefficient, and noise- or disturbance-affected modes, is introduced. This representation lays the groundwork for designing an effective FTC; 2) although few studies have focused on designing FTC strategies for both sensor and actuator faults in other applications (e.g., [15, 16, 17, 18, 19, 20]), most existing FTC schemes addressed only sensor faults [21, 6, 4] or actuator faults [22, 23, 24]. However, the MFHCA framework is the first FTC method designed specifically for hydraulic-powered HD-WMR systems, capable of handling both actuator and sensor faults; it features lower implementation complexity and requires only four design parameters for tuning; 3) the MFHCA framework remains entirely independent of any information from the modeling system, and it achieves strong control stability with exponential convergence rate.
The rest of the paper is organized as follows: after a brief explanation of the mechanisms of the studied hydraulic-powered HD-WMR, Section II presents the inverse kinematics of HD-WMRs to map the commanded motion in the robot’s base frame to the four driving speeds and rear/front steering angles of independent wheels. Through an in-depth exploration of the modeling of different sensor and actuator faulty modes, new representations of the motion dynamics of HD-WMRs are given in Section III. The step-by-step MFHCA framework design and a detailed stability analysis are given in Section IV. Section V presents experimental fault scenarios exposed to a 6,500-kg HD-WMR under rough and snowy terrains to evaluate the robustness and responsiveness of the MFHCA framework.
In this paper, $\mathbb{R}$ is used to denote the set of real numbers, while $\mathbb{R}^{n}$ refers to an $n$-dimensional real vector space. The symbol $(\cdot)^{\top}$ represents the transpose operation, and $\lambda_{\max}(\cdot)$ and $\lambda_{\min}(\cdot)$ correspond to the largest and smallest eigenvalues, respectively. The notation $\mathbb{R}^{\ell}\rightarrow\mathbb{R}^{n}$ denotes a function that is defined from a $\ell$-dimensional real space $\left(\mathbb{R}^{\ell}\right)$ to a $n$-dimensional real space $\left(\mathbb{R}^{n}\right)$. Define ${}^{b}\boldsymbol{x}_{j}^{w}$ as the variable $\bm{x}$ of the frame $j$ relative to the frame $b$, in the frame $w$.
@@ V-B Implementation of the MFHCA Framework
To investigate the robustness and tolerance of the MFHCA framework in the studied HD-WMR, we introduced artificial sensor and actuator fault scenarios for all four wheels at different time intervals, as illustrated in Figs. 4 and 5. The input of the MFHCA framework includes: 1) the errors between the faulty linear velocities $v_{i}$, derived from the faulty magnet-based speed sensor, and the command linear velocities in the wheel frame $v_{c,i}$, being inverse kinematically mapped from the reference velocities $v_{c}$ and $\omega_{c}$ in the WMR’s base frame; 2) the errors between the steering angles $\phi_{i}$, derived from wire-based encoders, and the reference angular velocities in the wheel frame $\phi_{c,i}$, being inverse kinematically mapped from $v_{c}$ and $\omega_{c}$. Finally, the MFHCA framework generates appropriate power efforts $u_{a,i}$ and $u_{s,i}$ in independent valve-regulated wheels. The parameters of the MFHCA framework used in this study are as follows: $\lambda_{i}=1,\xi_{i}=1$, $\beta_{i}=0.001$, and $k_{p,s}=1$. Communication among components of test-case HD-WMR was established in the Beckhoff IPC CX2030 before the operation (see Fig. 6).
The duty cycle of the experiment, conducted under the fault scenarios, is presented in Fig. 7. As indicated, the duty cycle was carried out on icy and rough terrain at $-8^{\circ}\mathrm{C}$, involving backward and forward movements, including the ascent and descent of two steep slopes from both the front and rear sides. The maximum commanded rotation of the HD-WMR, resulting from the joystick commands $\phi_{c,i}$, was $\theta_{b}=35$ degrees. Fig. 8 shows the adaptive parameters $\hat{\Psi}_{i}$ during the duty cycle for detecting and isolating faults. The second wheel (FR) experienced a significant jump due to the immediate stuck failures at $t=30$, $100$, and $200$. The results indicate that most short-term compensation efforts were directed at the second wheel, while most long-term efforts focused on the third wheel, primarily due to disturbances caused by ice slippage and the combined effects of sensor and actuator faults. Fig. 8 illustrates fault detection between $t=50$ and $t=235$, plus the immediate stuck failure in the sensor of the FR wheel at $t=30$, adhering to the fault scenarios indicated in Figs. 4 and 5. The tracking of linear velocities for the four wheels is shown in Fig. 9, illustrating how the wheels adhered to the commands even under fault conditions.
@@
Fig. 10 illustrates the four driving valve control signals $u_{a,i}$, derived from the output of Eq. (14), to force velocities of wheels $v_{a,i}$ to track the reference linear velocities $v_{c,i}$ while compensating for fault effects. The change in the sign of these signals indicates a change in the direction of the HD-WMR and highlights that the majority of efforts occurred during fault periods. In addition, Fig. 11 illustrates the steering valve-regulated control proposed in Eq. (15).
@@
The input is the information of the wire encoders, which show the positions of the two front and rear steering valve cylinders (converted into wheel angles $\phi_{i}$), and the output includes the proposed steering tracking control signals $u_{s,i}$ to track joy-stick-based commands $\phi_{c,i}$.
Table IV presents the comparative performance of state-of-the-art controllers for HD-WMRs, validating the capabilities of the MFHCA framework with fewer design parameters. It demonstrates better average tracking error and reduced valve control effort across all four wheels under identical conditions and snowy terrains. In addition, the implementation complexity is lower due to the reduced number of control design parameters that require tuning. Note: ‘avg.’ refers to average, ‘no.’ to number, ‘trk. err.’ to tracking error, ‘ctrl. eff.’ to control effort, and ‘des. params’ to the number of the required design parameters.
@@ VI Conclusion
Using Young’s inequality:
@@
This means that the Euclidean norm of the tracking linear velocity error vector exponentially converges to a stable region [32]. The radius of this region is given by $\sqrt{2\left(\lambda_{\min}(\boldsymbol{a})\right)^{-1}\sigma\eta^{-1}}$. This implies that the stability region will be larger if the intensity of faults, represented by $\sigma=\sum_{i=1}^{4}\frac{\left(F_{i}^{*}+D_{i}^{*}+\lambda_{i}\xi_{i}\right)^{2}}{2\sigma_{i}}$, is greater.

### v2

@@ I Introduction
Both light-duty and heavy-duty wheeled mobile robots (HD-WMRs) can to perform tasks that are inefficient, unsafe, or unfeasible for humans. However, while light-duty WMRs are typically used in structured environments, like warehouses and laboratories, HD-WMRs excel in challenging applications in rough terrains and under heavy-load conditions [1]. Many HD-WMRs are increasingly equipped with in-wheel hydraulic drive systems, each actuated by its motor and paired with front and rear steering mechanisms to improve maneuverability under heavy loads, enhance responsiveness to off-road conditions, and enable independent power delivery [2].
@@
To address the mentioned challenges, this paper introduces a novel model-free hierarchical control with fault accommodation (MFHCA) framework for highly complex and multi-stage HD-WMRs. It aims to accommodate both sensor and actuator faults passively in hydraulically powered HD-WMRs with independently controlled wheels, while ensuring continuous operation. Contributions of this work are summarized, as follows: 1) a new mathematical representation of the motion dynamics of HD-WMRs, incorporating various sensor and actuator modes, such as healthy, stuck failure, no signal, inefficient, and noise- or disturbance-affected modes, is introduced. This representation lays the groundwork for designing an effective FTC; 2) although few studies have focused on designing FTC strategies for both sensor and actuator faults in other applications (e.g., [15, 16, 17, 18, 19]), most existing FTC schemes addressed only sensor faults [20, 6, 4] or actuator faults [21, 22, 23, 24]. However, the MFHCA framework is the first FTC method designed specifically for hydraulic-powered HD-WMR systems, capable of handling both actuator and sensor faults; it features lower implementation complexity and requires only four design parameters for tuning; 3) the MFHCA framework remains entirely independent of any information from the modeling system, and it achieves strong control stability with exponential convergence rate.
The rest of the paper is organized as follows: after a brief explanation of the mechanisms of the studied hydraulic-powered HD-WMR, Section II presents the inverse kinematics of HD-WMRs to map the commanded motion in the robot’s base frame to the four driving speeds and rear/front steering angles of independent wheels. Through an in-depth exploration of the modeling of different sensor and actuator faulty modes, new representations of the motion dynamics of HD-WMRs are given in Section III. The step-by-step MFHCA framework design and a detailed stability analysis are given in Section IV. Section V presents experimental fault scenarios exposed to a 6,500-kg HD-WMR under rough and snowy terrains to evaluate the robustness and responsiveness of the MFHCA framework. In this paper, $\mathbb{R}$ is used to denote the set of real numbers, while $\mathbb{R}^{n}$ refers to an $n$-dimensional real vector space. The symbol $(\cdot)^{\top}$ represents the transpose operation, and $\lambda_{\max}(\cdot)$ and $\lambda_{\min}(\cdot)$ correspond to the largest and smallest eigenvalues, respectively. The notation $\mathbb{R}^{\ell}\rightarrow\mathbb{R}^{n}$ denotes a function that is defined from a $\ell$-dimensional real space $\left(\mathbb{R}^{\ell}\right)$ to a $n$-dimensional real space $\left(\mathbb{R}^{n}\right)$. Define ${}^{b}\boldsymbol{x}_{j}^{w}$ as the variable $\bm{x}$ of the frame $j$ relative to the frame $b$, in the frame $w$.
@@ V-B Implementation of the MFHCA Framework
To investigate the robustness and tolerance of the MFHCA framework in the studied HD-WMR, we introduced artificial sensor and actuator fault scenarios for all four wheels at different time intervals, as illustrated in Figs. 4 and 5. The input of the MFHCA framework includes: 1) the errors between the faulty linear velocities $v_{i}$, derived from the faulty magnet-based speed sensor, and the command linear velocities in the wheel frame $v_{c,i}$, being inverse kinematically mapped from the reference velocities $v_{c}$ and $\omega_{c}$ in the WMR’s base frame; 2) the errors between the steering angles $\phi_{i}$, derived from wire-based encoders, and the reference angular velocities in the wheel frame $\phi_{c,i}$, being inverse kinematically mapped from $v_{c}$ and $\omega_{c}$. Finally, the MFHCA framework generates appropriate power efforts $u_{a,i}$ and $u_{s,i}$ in independent valve-regulated wheels. The parameters of the MFHCA framework used in this study are as follows: $\lambda_{i}=1,\xi_{i}=1$, $\beta_{i}=0.001$, and $k_{p,s}=1$. Communication among components of test-case HD-WMR was established in the Beckhoff IPC CX2030 before the operation (see Fig. 6). The duty cycle of the experiment, conducted under the fault scenarios, is presented in Fig. 7. As indicated, the duty cycle was carried out on icy and rough terrain at $-8^{\circ}\mathrm{C}$, involving backward and forward movements, including the ascent and descent of two steep slopes from both the front and rear sides. The maximum commanded rotation of the HD-WMR, resulting from the joystick commands $\phi_{c,i}$, was $\theta_{b}=35$ degrees. Fig. 8 shows the adaptive parameters $\hat{\Psi}_{i}$ during the duty cycle for detecting and isolating faults. The wheel (FR) experienced a significant jump due to the immediate stuck failures at $t=30$, $100$, and $200$. The results indicate that most short-term compensation efforts were directed at the second wheel, while most long-term efforts focused on the third wheel, due to disturbances caused by ice slippage and the combined effects of sensor and actuator faults.
@@
Fig. 8 illustrates fault detection between $t=50$ and $t=235$, plus the immediate stuck failure in the sensor of the FR wheel at $t=30$, adhering to the fault scenarios indicated in Figs. 4 and 5. The tracking of linear velocities for the four wheels is shown in Fig. 9, illustrating how the wheels adhered to the commands even under fault conditions. Fig. 10 illustrates the four driving valve control signals $u_{a,i}$, derived from the output of Eq. (14), to force velocities of wheels $v_{a,i}$ to track the reference linear velocities $v_{c,i}$ while compensating for fault effects.
@@
The change in the sign of these signals indicates a change in the direction of the HD-WMR and highlights that the majority of efforts occurred during fault periods. In addition, Fig. 11 illustrates the steering valve-regulated control proposed in Eq. (15). The input is the information of the wire encoders, which show the positions of the two front and rear steering valve cylinders (converted into wheel angles $\phi_{i}$), and the output includes the proposed steering tracking control signals $u_{s,i}$ to track joy-stick-based commands $\phi_{c,i}$. Table IV presents the comparative performance of state-of-the-art controllers for HD-WMRs, validating the capabilities of the MFHCA framework with fewer design parameters. It demonstrates better average tracking error and reduced valve control effort across all four wheels under identical conditions and snowy terrains. In addition, the implementation complexity is lower due to the reduced number of control design parameters that require tuning. Note: ‘avg.’ refers to average, ‘no.’ to number, ‘trk. err.’ to tracking error, ‘ctrl. eff.’ to control effort, and ‘des. params’ to the number of the required design parameters.
@@ VI Conclusion
Using Young’s inequality [32, 33]:
@@
This means that the Euclidean norm of the tracking linear velocity error vector exponentially converges to a stable region [34, 35]. The radius of this region is given by $\sqrt{2\left(\lambda_{\min}(\boldsymbol{a})\right)^{-1}\sigma\eta^{-1}}$. This implies that the stability region will be larger if the intensity of faults, represented by $\sigma=\sum_{i=1}^{4}\frac{\left(F_{i}^{*}+D_{i}^{*}+\lambda_{i}\xi_{i}\right)^{2}}{2\sigma_{i}}$, is greater.